OpenStudy (anonymous):

Solve the following system of equations. x + y = 14 -x + y = 16

OpenStudy (anonymous):

-x + y = 16 y = 16+x subsitute this equation to x + y = 14 , then x + y = 14 x + (16+x) = 14 2x + 16 = 14 2x = 14 - 16 2x = -2 x = -2/2 x = -1

OpenStudy (anonymous):

Would it be (-1, 15)?

OpenStudy (anonymous):

@gerryliyana

OpenStudy (anonymous):

no no nevermind @sfregozo

OpenStudy (anonymous):

so its not -1,15 then?